OpenStudy (anonymous):

Use the table of the function f(x) = x4 - 4x to answer this question: x f(x) -2 24 -1 5 0 0 1 -3 2 8 What is the average rate of change from x = 0 to x = 2?

myininaya (myininaya):

The slope formula is change of y over change of x. So you say you don't recall this formula?

OpenStudy (anonymous):

oh ok i recognize that formula. im going to do that real quick

OpenStudy (johnweldon1993):

The average rate of change will be \[\large \frac{f(b) - f(a)}{b - a}\] a being our lower bound and b being the higher bound...so from 0 to 2 means a = 0 b = 2 f(b) means what the equation comes out to when you plug in 'b' for 'x' (or plug in 2 into the equation for 'x') f(a) is the same but for when you plug in 0 for 'x'

OpenStudy (johnweldon1993):

so f(b) what does the equation equal when x = 2? (in your table provided, it looks like 8) and f(a) what does it equal when x = 0? (again in the table, it comes out to 0) so here \[\large \frac{f(b) - f(a)}{b - a} = \frac{8 - 0}{2 - 0} = \frac{8}{2} = ?\]

OpenStudy (anonymous):

ok sorry i took so long i was trying understand your work as best as possible. so the final answer is 4. thanks you so much for your help i understand it now thank you!

OpenStudy (johnweldon1993):

No problem at all, is there anything you need clarification on? :)

OpenStudy (anonymous):

nope you explained it verry well (:

OpenStudy (johnweldon1993):

Very good :)

OpenStudy (anonymous):

(;

OpenStudy (johnweldon1993):

;P